OpenStudy (anonymous):

(2x+5) (3x-4) (6x-7)=0 solve plss..

OpenStudy (anonymous):

x = -5/2, x = 4/3, x = 7/6

OpenStudy (anonymous):

got it ?

OpenStudy (anonymous):

yes thanks